A Complete Bibliography of Publications in Computers and Mathematics with Applications (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 1.03

Title word cross-reference

(4, 5) [SH77]. 0 [JS78a]. 1 [JS78a]. $10.95 [Cho77a]. $11.25 [Pol76]. $12.50 [Ano77e]. $15.95 [Kot79]. $16.95 [Ano77d, Ano77h]. $17.50 [Orr79a]. $18.00 [Kin77]. $18.50 [Wal78]. $18.95 [Ano77b, Coo78a]. $19.50 [Ste78e]. $19.95 [Cho77b]. $22 [Ell76]. $22.50 [Rod77b]. $22.95 [Orr79b, Ste78d]. $23 [SS76]. $24.95 [Ano77a, KS78c, Rod77a]. $25.00 [Orn77, Ano77f, Orn77]. $27.50 [Ste78b]. $28.95 [Ano77g]. $29.95 [Ano77c]. $32.90 [Ste78c]. A [EH78, LP78]. C^{1} [MM79a]. f [Mil75]. f^{∗}(0)f − f(0)f^{∗} [Mil75]. k [AEMT79]. λ [Tur75]. n [ES79]. p [AH79]. Q [Nah76]. x [Wai79].

-cube [ES79]. -function [Nah76].

-parametric [AH79]. -projection [Wai79]. -stable [LP78]. -th [AEMT79].

11 [Ste78e]. 1977 [Zei79a, Zei79b, Zei79d]. 1979 [Ano78d]. 1980 [Ano79c].

236pp. [Wid79].

99pp [Pol76].

a-posteriori} [GTG75]. Academic

[Rod77b]. accelerated [LI75]. accelerating [Smi78]. Acceleration [Wai79]. Adams [KFS77, RT78]. Adams-type [RT78]. addressing [Dym79]. Adleman [BB79]. Advanced [Ort75b]. Advantages [Mit79]. Algebra [Zei79a]. algorithm [CC76, Coo79, Fat76, GM76, JS78a, JS78b, KS75a, Nah76, Rie76, Wei77]. Algorithmic [Cha76]. Algorithms [KMS77, WK77, GK76, GPV78, Ste78c, Coo78a]. alternating [AH79]. always [BB79]. amount [Str76]. Analysis [Ano77c, Cho77b, Cho78b, GPV78, Rod77a, Wid79, Ano79c, BJY78, Bro79, CO75, Cho78a, LG75, Nah76, Sto79, Sza79, War75, Ano77b, Ano77e, Ano77g, Wid78]. Analytical [Far77]. Andrews [Zei79c]. Anthony [KS78c]. Anthropomorphic [SS76]. antiparticles [Fra77]. Application [Del78, Asa76, Zei79a]. Application-Oriented [Zei79a]. Applications [Ng77, HP75, She78, Coo77b]. Applied [Kot79, Ano79c, Gre77, Ano77b]. approach [LeB77, LA79, MTC75]. approaches [KS78a]. approximating [Mit79]. approximation [AEMT79, Ara75, FN77, GM76, MM79a, Net78, Ost75]. approximations [Luk77, Ort75a, RO77]. arbitrary [KPR78]. Archaeology [Kin77]. Arithmetic [Gre77, MT77]. aspects [HP75, Str76]. Asymptotic [LR75, Won77, Luk75]. August [Ano77b, Ano77c, Ano77h, Ano78d]. automated [OMW76]. Automatic [Rua79, Win76]. Avriel [Rod77a]. axisymmetric [KZ75].

B [Ano77f, Wal78]. based [LFW75]. bases [AM79, McL79c]. basis [Wac79]. Bathe [Ano77g]. be [GK76]. Beach [Zei79d]. beam [KZ78, KS78b]. behavior [LR75, Str76]. Belgrade [SS76]. bending [Pea76]. Berlin [Ste78c]. Bessel [Luk75]. Better [Sha79]. Bicubic [Dim77]. biharmonic [Red77]. Biology [Ell76]. Bird [Ano77h]. Blaauw [Ano77d]. board [Ano76a, Ano77i, Ano78b, Ano75a, Ano79a]. Bobiller [Wal78]. Book [Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77a, Ano77g, Ano77h, Cho77a, Cho77b, Cho78a, Cho78b, Coo77b, Coo78a, Coo78b, Ell76, Kin77, KS78c, Kot79, Orn77, Orr79b, Orr79a, Pol76, Rod77b, Rod77a, Ste78d, Ste78a, Ste78e, Ste78b, Ste78c, SS76, Wal78, Wid78, Wid79, Zei79a, Zei79b, Zei79d, Zei79c, Ano76b]. Books [Ano78a, Ano77k, Ano79b]. both [KZ78]. boundaries [MM79a, Mit79]. boundary [FG77, LL76, RO77, Sac76, She78, dVDB75]. boundary-value [RO77, She78]. bounded [MTC75]. bounds [Dim77, GTG75]. Boyd [ZE77]. buckling [KSZ75, KS78b]. built [KZ78]. Burger [Sac76]. Butterworth [Zei79c].

C [Cho77a, Kot79, Wal78, Wid78].

calculation [EH78, SS78]. calculations [Wei77]. calculator [HS77]. Cambridge [Kin77]. Canada [Ano79c]. candidate [GF79]. capacity [Wei77]. carrying [MT77]. certain [JJ77]. characteristic [Rao78]. characteristics [LVW76]. charge [Coo75, RC76]. checkers [DS79]. Cheung [Ano77e]. Choi [Kot79]. Ciriani [Ste78b]. class [CS77, FT75, LP78, Sac76, Wai79]. classic [KMS77]. Classifications [Wid79]. Cliffs [Ano77g, Coo78a, Kot79, Orr79a, Wal78, Wid78, Zei79b]. coefficients [Win79]. combination [Tal78, Tal79]. combinations [DM77]. combinatorial [Ste78a, Coo78a]. Communications [Pol76]. Comparing [Gar76]. comparison [MM79b]. Compilation [Ano77f]. Complete [Win76, KPR78, Sil79]. Complexity [OMW76, Orr79b, Ste78d]. Computation [HS77, Rho77, Dym79, Rao78]. Computational [HP75, JS78a, Neu75, Ano77c]. Computations [Ng77, CK76]. Computer [Cho77a, Coo77b, Orr79a, Zei79b, Ara75, HP75, Knu75, SS78, Ste78e, Sto79]. Computers [Kin77, Nie75, GPV78, KS78c, Pol76].

Computing [Orr79b, Ste78d]. concave [CC76]. conceal [BB79]. concept [Del78, War76]. concerning [DGN75]. conditioning [Del78]. conditions [FG77, MR76]. configurations [KS78b]. confirmation [ZE77]. conforming [KPR78, Pea76]. connection [Sze75]. consistence [Tho78]. constant [Knu76]. constrained [DK76]. Construction [Gou79a]. Continued [FD78]. control [BJY78, MR76, MTC75]. convergence [Pet78, Sim78, Smi78, Tal78, Tal79]. conversion [Asa76]. coolant [CU78]. Cooper [Ste78e]. Cornelius [Hor77, McC75, RBE^{+}75]. Corrector [RT77, LFW75]. correctors [KFS77]. cosine [KS75a]. Cosmology [Sze75]. coupled [JL79]. criterion [DM77, Tal79]. Critical [CU78]. Cross [Wid79]. Cross-Classifications [Wid79]. cryptosystems [BB79]. cube [ES79]. Cubic [JL79]. cubics [Mit79]. curvature [DGN75]. curved [AM79, MM79a, McL79c, Wac79]. cybernetics [Zei79d].

D [Rod77b]. Data [Ano77f, Cho78b, Cho78a, JJ77]. dates [Rho77]. David [Orr79a, Zei79b]. Davidenko [KZHH77]. Day [Cho77a]. defined [Luk77]. Definition [War76]. deflections [KZ78]. degree [Gou79b]. demodulation [Win76]. dense [EGS75]. density [FG76, LR75]. Deo [Coo78a]. design [LA79, SS78]. designing [CO75]. Detecting [SH77]. determination [Coo77a]. developments [Sza79]. dielectrics [RC77]. difference [Ada75, Ros75]. Differential [BKLV75a, BKLV75b, Bro79, Far77, Fat76, Fat78, GTG75, LP78, Rie76, Ros78, Win79]. Diffusion [Ara75, FN77]. digital [KS78c, Ano77d]. dimensional [McL79c]. direction [AH79]. discrete [LeB77, Wac77, Orr79a, Zei79b]. discriminant [CO75]. dissection [GPV78]. Dissemination [Ano77f]. distance [FG76]. distillation [LL76]. distributions [Cho77c]. divergence [DM77]. division [Asa76]. do [BB79]. Donald [Orr79a, Zei79b]. Donelley [Zei79a]. Doran [Kin77]. Downstream [FG77]. Dreyfus [Ano77f]. dual [GM76, KS75a, KMS77, LA79]. Dublin [McC75]. Dun [Zei79a]. Dunkelberg [Wid78]. dynamic [FT75].

Early [Ort75b]. easy [BKLV75b]. Edited [KS78c, Ste78b]. Editorial [Ano76a, Ano77i, Ano78b, Ano75a, Ano79a]. education [Nie75]. effect [JJ77]. efficient [LA79, Wei77]. eigensystems [Ger77]. eigenvalue [KM78]. eigenvalues [EH78, KS78a]. Eiselt [Ste78c]. elastic [KZ75, KZ78]. elasticity [Pea76]. electric [Wei77]. Element [Ano77b, Ano77g, And79, DN78, FN77, FGN79, GM76, Gou79a, Gou79b, Mit79, Net78, Pet78, Red77, RO77, Rua79, She78, Sza79, Tho78, War75]. elements [AM79, KPR78, McL79c, Pea76, Wac79, Ano77a]. Elucidation [DGN75]. ends [KZ78]. engineering [Ano79c]. England [Ano77e]. Englewood [Ano77g, Coo78a, Kot79, Orr79a, Wal78, Wid78, Zei79b]. enhancements [OMW76]. Enslein [KS78c]. entropy [Cha76]. enumeration [JS78b]. Epley [ZE77]. equation [FD78, Red77, Rie76, Ros75, Sac76, Rod77b]. equations [Ada75, DN78, Far77, Fat76, Fat78, GTG75, GK76, JL79, KMS77, LP78, Ros78, Wac77, WK77, Wai79, Win79, dVDB75]. Erratum [Ano78c]. Error [Rao78, AAG79, GTG75, Gou79b, Luk77, Nah76, RO77, Tho78]. Error-free [Rao78]. estimate [LR75, Ost75]. estimates [AAG79, Gou79b, RO77]. Estimation [KZ78, LVW76]. estimator [LM75]. evaluating [CO75]. Evaluation [And79, AMT78, AEMT79, Ano77f, Knu76]. Evolution [ES79]. exact [CC76, Sac76]. example [MP76]. examples [Ste78e]. Existence [Red77]. expansions [Luk75, Won77]. experience [JS78a]. extended [Tho78]. extensibility [DS79]. Extensions [McL79a]. extrapolated [AH79].

F [Cho77a, Kin77, Orr79a, Zei79b, Zei79d]. Factorisation [EO79]. faculty [Coo77a]. family [CS77]. fault [FG76]. Feature [DM77]. Fehlberg [SH77]. fields [War75]. Finite [Ano77b, Ano77a, Ano77g, FN77, Net78, Pet78, War75, Ada75, And79, AM79, DN78, FGN79, GM76, Gou79a, Gou79b, KPR78, McL79c, Pea76, Red77, RO77, Rua79, She78, Sza79, Tho78, Ano77e]. fire [Cek78]. first [Rie76]. Fisher [Zei79a]. fitting [MP76]. fixed [Cho77c, Coo75, Ost75, Pet78, RC76]. flow [Bal75, CU78, CK76, FG77]. forests [Cek78]. Foreword [McL79b, Rod75]. Formalization [ZE77]. formula [Asa76]. formulae [Sil79]. formulas [Cla76, LFW75, SH77]. formulation [AH79]. forthcoming [Ano75b, Ano75c, Ano76c]. foundation [KZ78]. foundations [Zei79d]. Fourier [Bat78, LG75]. fraction [FD78]. Frajer [Ste78c]. Francisco [Cho77a]. Free [Wid79, Rao78]. fronts [Cek78]. fuel [CU78]. function [JJ77, LM75, LR75, Nah76]. Functions [AMT78, AEMT79, Luk75, Luk77, Sim78, Wac79, Won77]. fuzzy [KS78a].

G [Cho77b, Cho78a, Cho78b, Ste78a, Zei79c]. game [Fra77]. games [FT75]. gamma [Mar77]. general [AH79]. generalisations [McL79a]. Generalised [EO79]. generalized [HP75]. generated [Bro79]. generating [Mar77, Wei77]. generation [Rua79, Ano77f]. Geography [Ano77c].

George [Zei79d]. Gerrit [Ano77d]. Gilb [Coo78b]. global [MP76]. Gnanadesikan [Cho78a, Cho78b]. Gordon [Zei79d]. GPSS [Wal78]. graphs [CK76, EGS75]. grid [JJ77, Pet78]. Gruyter [Ste78c].

H [Wid79, Zei79d]. Hall

[Ano77d, Ano77g, Coo78a, Kot79, Orr79a, Rod77a, Wal78, Wid78, Zei79b]. hand [HS77]. handbook [Ste78c]. Hankel [Won77]. hard [BKLV75b]. Harvard [Kin77]. having [CU78]. Heat [Rod77b]. Hebrew [Rho77]. Helmut [Ste78c]. Henschke [Cho77a]. Herbert [KS78c, Kea75]. hereditary [MR76]. Hermite [Cla76, Gou79a, Gou79b]. Hermitian [Ada75]. heuristic [Coo75, CS77, LeB77]. Hierarchies [Pea76]. high [McL79a, MM79b, McL79c, Red79]. higher [AM79]. Hodson [Kin77]. Holden [Cho77a]. Holden-Day [Cho77a]. Horst [Ste78c]. Hybrid [LI75, MTC75]. hydrology [Ste78b]. hyper [Ove75]. hyper-resolution [Ove75]. Hypergeometric [Ng77].

identification [Del78]. IEP [Zei79a]. II [Coo77b]. imbedding [KSZ75, KM78, KS78b, LM75, LVW76]. implementation [Ove75, Ano77d]. implicit [AH79, JS78b]. improved [Coo79]. increased [RT77, RT78]. Indirect [Dym79]. infinite [HS77]. initial [KZM78]. Innovative [Ano79c]. Institute [SS76, Ort75b]. integer [CC75, Coo79, JS78a, JS78b]. integral [BKLV75b, DN78, FD78, MR77]. integrals [And79, HS77, Luk77]. Integration [AM79, dVDB75]. International [Ano78d, Ano79c]. interpolants [LFW75]. Interpolation [Gou79b, AAG79, Cla76, JJ77, McL79c]. Interscience [Cho77b, Cho78a, Cho78b, Coo77b, Ell76, KS78c, Ste78d, Ste78e, Ste78b]. Introduction [Ano77a, KS75b, Cho77a, Ell76]. Introductory [Kot79]. Invariant [LVW76, LM75]. inversion [HP75, MR77]. isoparametric [And79, AM79, MM79b]. issues [Ano75b, Ano75c, Ano76c]. Iterations [Gar79].

J [Ano77b, Ano77a, Coo78a, Kin77, Ste78b, Zei79c]. James [Ste78e, Zei79a]. Jersey [Ano77g, Kot79, Zei79b]. John [Ano77b, Ano77c, Ano77a, Ano77h, Orr79b, Ste78d, Ste78a, Wid78]. July [Rod77a]. June [Ano77d, Ano77g, Ano79c].

Kahan [Wal78]. Katta [Ste78a]. key [BB79]. Kleinrock [Coo77b]. Kossack [Cho77a]. Kurt [KS78c].

L [Zei79a]. laboratory [Ste78e]. Laguerre [Gar76, Hai78]. laminar [dVDB75]. Lanczos [But75, DGN75, Hor77, McC75, Ort75b, Tau75, RBE^{+}75]. language [KS75b]. languages [War76]. large [KZ78, LeB77]. Larry [Ano77b]. latent [Tur75]. Laver [Pol76]. layer [KZ75, dVDB75]. least [FGN79]. Legend [But75]. Legged [SS76]. Leonard [Coo77b]. Lieberstein [Kea75]. Linear [Ng77, CC75, Coo79, JL79, LI75, RC77, She78, Tal78, Tal79, WK77, Wai79, Win79, Ste78a]. linearly [DK76]. Liouville [Bal75]. loads [KSZ75]. local [DGN75]. location [Coo78c, LeB77]. Locomotion [SS76]. logarithmic [Won77]. London [Zei79d, Zei79c]. long [EGS75].

M [Ano77c, Coo78a, SS76]. Machines [Ano77h]. magnetic [War75]. Maione [Ste78b]. majority [GF79]. Marcum [Nah76]. Massachusetts [Kin77]. Mathematical [Ano77a, Ano78d, Ell76, Lie75, Ste78b, Zei79c]. Mathematics [Orr79a, Zei79b, Gre77, Nie75, Orn77, Kin77]. Mather [Ano77c]. matrices [KS78a, Tur75]. matrix [Del78, EH78, HP75, KM78, Rao78]. Maximization [DK76]. McAllister [Orr79a, Zei79b]. McLone [Zei79c]. Mechanisms [SS76]. Melvin [Kea75]. memoriam [Hor77, Kea75]. memory [FG76, RBE^{+}75]. meshes [Rua79]. messages [BB79]. Metamathematical [DS79]. Method [Ano77e, EO79, Gar76, Ada75, Coo75, DN78, KSZ75, KZHH77, KM78, KZM78, KS78b, LVW76, Mar77, MP76, Neu75, Ort75a, Red79, Ros78, Smi78, Tho78]. Methods [Ano77c, Ano77g, BKLV75b, CS77, FGN79, GK76, Ger77, JJ77, KS78c, LP78, McL79a, MM79b, RT77, RT78, She78, Ste78c, WK77, Wai79, Rod77a, Cho78a, Cho78b]. metrics [Coo78b]. Microcomputers [AMT78, AEMT79]. Mihailo [SS76]. mini [Ste78e]. mini-computer [Ste78e]. minimization [JJ77]. mixed [She78]. model [Knu75, Red77, ZE77]. Modeling [Zei79c]. models [CO75, Dym79, Ste78b]. modern [Orn77]. modification [KZHH77]. monomial [Mil75]. Montréal [Ano78d, Ano79c]. Mordecal [Rod77a]. Moulton [KFS77]. multicomponent [LL76]. Multiple [Cho77b]. multiplication [Asa76]. multiprogrammed [Ara75]. multistep [LP78]. Multivariate [Ano77c, Cho78b, Cho78a]. Murray [Pol76]. Murty [Ste78a].

N [Ano77a, Coo78a]. N.J [Orr79a]. necessary [MR76]. network [LA79]. Newton [Gar76, Red79]. Nievergelt [Coo78a]. Nine [Ros75]. Nine-point [Ros75]. NJ [Coo78a, Wal78, Wid78]. Nodal [KPR78]. node [And79]. noisy [Win79]. Non [CC75, RC77, Coo79, DGN75, JL79]. Non-linear [CC75, RC77, Coo79, JL79]. non-local [DGN75]. Nonlinear [LL76, WW78, Bro79, CS77, FD78, FN77, GM76, GTG75, GK76, KZHH77, LP78, LI75, Net78, RO77, War75, Rod77a]. normal [Cho77c]. note [RO77, Sim78]. notice [Ano77j]. nozzle [Bal75]. Numerical [BJY78, Gar79, Ger77, KS78a, MR76, MR77, SS78, Ano79c, Fat76, Fat78, LP78, MM79b, Wac77, dVDB75, Ano77g]. numerically [Bro79].

O [Cho77b]. Obituary [Ano77j]. Observations [Cho78b, Cho78a]. Oden [Ano77a]. ODEs [Sha79]. one [Cho77c, Fat78, Smi78]. one-parameter [Smi78]. one-step [Fat78]. operating [ZE77]. Operations [Ste78c]. Operator [Ros78, Ost75]. opinion [Wid78]. Optimal [LM75, Coo77a, MTC75, RO77, Tal78, Tal79]. Optimally [KFS77]. optimization [AH79, LI75]. order [AM79, Far77, Fat78, KPR78, McL79a, MM79b, McL79c, Red79, Rie76]. ordinary [Fat76, Fat78, GTG75, Ros78]. Oriented [Zei79a]. Orr [Ger77]. orthogonal [KZM78]. Other [Ano76b, Ano77k, Ano79b, Orn77]. outsiders [Orn77]. Oxford [Pol76].

P [Ano77c, Wal78]. Padé [Luk77]. page [FG76]. Papers [Ano75b, Ano75c, Ano76c]. parabolic [Ada75, JL79, MM79a, Net78, Wac77]. Parallel [Gar76, CU78, EH78, KFS77]. parameter [LVW76, Smi78]. parameters [KZ78]. parametric [AH79]. Part [Ort75a]. particles [Fra77]. parts [DGN75]. Passover [Rho77]. patch [Dim77]. paths [EGS75]. Paul [Orn77]. PDP [Ste78e]. PDP-11 [Ste78e]. penetration [KZ75]. Pergamon [Ano77e, Ano77f, Orn77]. Periodic [EO79]. perturbative [Ros78]. phases [Neu75]. Physical [Ano77c]. physicists [Orn77]. physiology [Lie75]. piecewise [MM79a, Ort75a]. plane [Pea76]. plate [Pea76]. plates [CU78]. point [Ost75, RO77, Ros75]. Poisson [Ros75]. polynomial [KPR78, Mil75, Ort75a, Rao78]. polynomials [DK76, KZM78, Rie76]. Porter [Knu76]. Post [KS78b]. posteriori} [GTG75]. power [Rie76]. pp [Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77a, Ano77g, Ano77h, Cho77a, Cho77b, Ell76, Kin77, KS78c, Kot79, Orn77, Orr79a, Rod77b, Rod77a, Ste78d, Ste78e, Ste78b, Ste78c, SS76, Wal78]. pp. [Cho78a, Cho78b, Coo77b, Ste78a, Wid78]. practice [Coo78a]. Predictor [LFW75]. predictors [KFS77]. Prentice [Ano77d, Ano77g, Coo78a, Kot79, Orr79a, Rod77a, Wal78, Wid78, Zei79b]. presence [WW78]. Press [Ano77e, Ano77f, Kin77, Orn77, Pol76, Rod77b, Wid79]. principles [Tho78]. probability [FG76, MT77]. problem [AH79, Bal75, Cla76, Coo75, CC76, Coo78c, Del78, FN77, Ger77, JS78b, LA79, Net78, RO77, RC76]. problems [BKLV75b, CS77, FG77, GM76, KM78, LeB77, LL76, Lie75, MR76, MTC75, Sac76, She78]. Probst [Wal78]. Procedures [Wid78, Tal78, Tal79]. processing [Wid78]. Processor [Tom75]. profiles [GF79]. Profitability [CK76]. program [CK76]. programmable [HS77]. Programming [Ano78d, Rod77a, CC75, CS77, Coo79, JS78a, JS78b, War76, Cho77a, Ste78a]. programs [OMW76, Ano77h]. projection [GK76, WK77, Wai79]. proof [DS79]. proof-checkers [DS79]. Propagation [Cek78]. properties [Sim78]. Proportions [GF79]. proving [OMW76, Win76]. public [BB79]. punch [KZ75]. Punctual [BJY78]. Pupin [SS76].

quadrature

[BKLV75a, BKLV75b, Sil79, WW78]. Queueing [Coo77b]. queues [Neu75].

R [Kin77, Ste78b, Wal78, Zei79c]. radix [Asa76]. Ralston [KS78c]. random [LG75]. ranges [RT77, RT78]. ratio [Rie76].

Rational [AAG79, Wac79, AEMT79, Cla76, Gou79a, Gou79b, LFW75, McL79a]. reactors [CU78]. real [EH78]. received [Ano76b, Ano77k, Ano78a, Ano79b]. recurrence [Cla76]. recursive [Asa76]. Reddy [Ano77a]. Regression [Cho77b]. regularized [Str76]. Reingold [Coo78a]. related [Cla76, OMW76]. relationships [Dym79]. Reliability [Win79]. remarks [Luk75]. replacements [LI75]. Representations [Hai78]. research [Ste78c, Wid78]. resolution [Ove75]. Review [Ano77b, Ano77c, Ano77d, Ano77e, Ano77f, Ano77a, Ano77g, Ano77h, Cho77a, Cho77b, Cho78a, Cho78b, Coo77b, Coo78a, Coo78b, Ell76, Kin77, KS78c, Kot79, Orn77, Orr79b, Orr79a, Pol76, Rod77b, Rod77a, Ste78d, Ste78a, Ste78e, Ste78b, Ste78c, SS76, Wal78, Wid78, Wid79, Zei79a, Zei79b, Zei79d, Zei79c]. Reynolds [Wid79]. Richard [Ano77h]. Riemann [DGN75, Tau75]. right [Rie76]. rigid [KZ75]. Rivest [BB79]. road [LA79]. Robots [SS76]. Roman [Orn77]. Root [AMT78]. Rooting [Gar79]. roots [AEMT79, Tur75]. routines [CO75]. Rubinow [Ell76]. rules [WW78].

S [Ell76, KS78c]. Saaty [KS78a]. sample [Cho77c]. San [Cho77a]. Savage [Orr79b, Ste78d]. scale [LeB77]. scheme [AH79, Fat78]. Science [Ano77f, Ano79c, Kot79, Orr79a, Zei79b]. second [Far77, Ano79c]. Segerlind [Ano77b]. selection [GK76]. semi [HS77, Wac77]. semi-discrete [Wac77]. semi-infinite [HS77]. semicardinal [Sil79]. sequences [Bro79, LG75, Smi78]. sequential [Cho77c, Dym79]. series [KS75a, Rie76, Smi78]. set [KS75b, War76]. SETL [KS75b]. sets [Cha76, KS78a]. Shamir [BB79]. shock [MP76]. side [AM79, Rie76]. Significance [MT77]. simulation [CO75, Wal78]. Simultaneous [Gar79]. singular [DN78]. singularities [Red79, Won77]. singularity [WW78]. Sixties [Ort75b]. size [JJ77]. sizes [Cho77c]. smoothing [Str76]. Society [Pol76]. software [Sha79, Coo78b]. Solution [EO79, Ada75, CS77, DN78, FD78, Fat76, Fat78, GM76, GTG75, MR76, Rie76, Ros78, Wac77]. solutions [Far77, Red77, Ros75, Sac76, Str76]. Solving [BKLV75b, GK76, KS75a, LP78, WK77, Wai79]. Some [Luk75, Str76, Sza79, Cla76, DGN75, Dym79, Lie75, Sim78, Orn77]. Sommerfeld [Ger77]. Sonquist [Wid78]. spacetime [DGN75]. Sparse [EO79, EGS75]. spectroscopy [Bat78]. spinor [Sze75]. spline [JL79, LR75, Sim78]. splines [BKLV75a]. splitting [Tau75]. Square [AMT78, Gar79]. Square-Rooting [Gar79]. squares [FGN79]. squeeze [Mar77]. Stability [Bro79, Gar79, Wac77, McL79c, RT77, RT78]. stable [KFS77, LP78]. Stack [FG76]. staffing [Coo77a]. Stanat [Orr79a, Zei79b]. Standard [Ste78c]. state [MTC75, Wei77]. stationary [Red77]. Statistical [Cho78b, Cho78a, KS78c]. Statistics [Cho77a, Kot79]. Step [Ort75a, Fat78]. Stieltjes [Luk77]. stiff [LP78]. stiffness [SH77]. stochastic [Coo78c, LeB77]. Strategy [FT75]. string [BJY78]. Strip [Ano77e]. Structural [Ano77e]. structured [Knu75]. Studies [Ort75b]. study [RC76, Tom75]. Sturm [Bal75]. Subspace [GK76]. successive [LI75]. Sung [Kot79]. surface [Ste78b]. Survey [Wid78]. Symbolic [Sto79]. Symposium [Ano78d, Ano79c]. System [Ano77d, Knu75, ZE77]. Systems [Coo77b, EO79, Ara75, FG76, GK76, KZHH77, LI75, WK77, Wai79, Wei77].

T [Ano77a, Ste78b, Wid79]. Tau [Ort75a]. technique [JL79, MM79a]. techniques [AM79, Wai79]. Technology [Ano77f]. temperatures [Hai78]. ten [And79]. ten-node [And79]. tensor [DGN75, Tau75]. Tenth [Ano78d]. test [Cho77c]. tetrahedral [And79]. th [AEMT79]. theorem [DS79, OMW76, Win76]. theorem-proving [OMW76]. theoretic [War76]. theoretical [KS75b]. theories [She78]. Theory [Ano77a, Coo78a, Neu75]. third [Gou79b]. thoughts [DGN75]. ties [FT75]. time [Dym79]. Tom [Coo78b]. transform [Bat78]. transformation [McL79a, MM79b]. Transformations [Ng77]. transforms [MR77, Won77]. transportation [CC76, Coo78c, LeB77, RC76]. transportation-location [Coo78c, LeB77]. treatment [Lie75]. triangular [Rua79]. Tridiagonal [EO79]. trigonometric [KMS77]. Two [Cho77c, McL79c, RO77, Rie76]. two-dimensional [McL79c]. two-point [RO77]. Two-sample [Cho77c]. type [FGN79, Gou79a, Gou79b, Mil75, RT78].

U [Ste78b]. unconstrained [JJ77]. uniform [Luk75]. uniqueness [Red77]. University [Kin77, Pol76, Coo77a]. urban [LA79]. used [GK76]. uses [Neu75]. using [MR76, Ste78e]. utilization [Tom75].

V [Rod77b, Wal78]. value [KZM78, LL76, RO77, Sac76, She78]. variable [Fat78]. variables [KPR78]. Variance [Cho77b]. variates [Mar77]. variational [GM76, Tho78]. varied [Str76]. vector [GPV78, Sto79, War76]. velocities [CU78]. verifiers [DS79]. Via [Gar79, GM76, KS78b]. vibrating [BJY78]. virtual [FG76]. Viscous [Bal75, FG77]. Volume [Coo77b, Orn77]. Vukobratovic [SS76].

W [Ste78e]. Wachspress [Gou79a, Gou79b, AAG79]. Wallis [Ste78b]. water [Ste78b]. weighting [LM75]. Wesolowsky [Cho77b]. where [Rie76]. Widder [Rod77b]. Wiley [Ano77b, Ano77c, Ano77a, Ano77h, Cho77b, Cho78a, Cho78b, Coo77b, Ell76, KS78c, Orr79b, Ste78d, Ste78a, Ste78e, Ste78b]. Wilf [KS78c]. William [Wid78]. Wilson [Ano77g].

xiv [Wid79]. xx [Coo77b].

Year [Rho77]. York [Ano77b, Ano77c, Ano77a, Ano77h, Cho77b, Cho78a, Cho78b, Coo77b, Ell76, KS78c, Orn77, Orr79b, Rod77b, Rod77a, Ste78d, Ste78a, Ste78e, Ste78b, Wid79, Zei79a].

References

Apprato:1979:RIW

[AAG79] D. Apprato, R. Arcangeli, and J. L. Gout. Rational interpolation of Wachspress error estimates. Computers and Mathematics with Applications, 5(4): 329–336, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                                  tronic).  URL http://www.
                                                  sciencedirect.com/science/
                                                  article/pii/0898122179900920.

Adam:1975:HFD

[Ada75] Yves Adam. A Hermitian finite difference method for the solution of parabolic equations. Computers and Mathematics with Applications, 1(3–4):393–406, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                                  sciencedirect.com/science/
                                                  article/pii/0898122175900413.

Andrews:1979:EFM

[AEMT79] M. Andrews, B. Eisenberg, S. F.

McCormick, and G. D. Taylor. Evaluation of functions on microcomputers: rational approximation of k-th roots. Computers and Mathematics with Applications, 5(3):163–167, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900385.

Avdelas:1979:GPF

[AH79] G. Avdelas and A. Hadjidimos.

On the general problem of formulation and optimization of a p-parametric extrapolated alternating direction implicit scheme. Computers and Mathematics with Applications, 5(1):51–57, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122181901395.

Andersen:1979:ITI

[AM79] C. M. Andersen and Robin J. Y.

McLeod. Integration techniques for isoparametric and higher order bases on finite elements with a curved side. Computers and Mathematics with Applications, 5 (4):285–295, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900890.

Andrews:1978:EFM

[AMT78] M. Andrews, S. F. McCormick, and G. D. Taylor. Evaluation of functions on microcomputers: Square root. Computers and Mathematics with Applications, 4 (4):359–367, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900160.

Andersen:1979:EIT

[And79] C. M. Andersen. Evaluation of integrals for a ten-node isoparametric tetrahedral finite element. Computers and Mathematics with Applications, 5(4):297–320, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122179900907.

Anonymous:1975:EB

[Ano75a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 1(1):ifc, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122175900012.

Anonymous:1975:PFIa

[Ano75b] Anonymous. Papers in forthcoming issues. Computers and Mathematics with Applications, 1(2):255, June 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900231.

Anonymous:1975:PFIb

[Ano75c] Anonymous. Papers in forthcoming issues. Computers and Mathematics with Applications, 1(3–4):428, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900450.

Anonymous:1976:EB

[Ano76a] Anonymous. Editorial board.

Computers and Mathematics with Applications, 2(1):ifc, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900018.

Anonymous:1976:OBR

[Ano76b] Anonymous. Other book received.

Computers and Mathematics with Applications, 2(2):191, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900328.

Anonymous:1976:PFI

[Ano76c] Anonymous. Papers in forthcoming issues. Computers and Mathematics with Applications, 2(1):99, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122176900109.

Anonymous:1977:BRI

[Ano77a] Anonymous. Book review: An Introduction to the Mathematical Theory of Finite Elements: by J. T. Oden and J. N. Reddy. John Wiley, New York, 1976. 419 pp. $24.95. Computers and Mathematics with Applications, 3(2):153, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812217790075X.

Anonymous:1977:BRA

[Ano77b] Anonymous. Book review: Applied Finite Element Analysis: by Larry J. Segerlind. John Wiley, New York, August 1976. 401 pp. $18.95. Computers and Mathematics with Applications, 3(2):154, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900803.

Anonymous:1977:BRC

[Ano77c] Anonymous. Book review: Computational Methods of Multivariate Analysis in Physical Geography: by P. M. Mather. John Wiley, New York, August 1976. 532 pp. $29.95. Computers and Mathematics with Applications, 3(2):153, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900736.

Anonymous:1977:BRD

[Ano77d] Anonymous. Book review: Digital System Implementation: by Gerrit A. Blaauw. Prentice–Hall, June 1976. 384 pp. $16.95. Computers and Mathematics with Applications, 3(2):153, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900748.

Anonymous:1977:BRF

[Ano77e] Anonymous. Book review: Finite Strip Method in Structural Analysis: by Y. K. Cheung. Pergamon Press, England, 1976. 228 pp. $12.50. Computers and Mathematics with Applications, 3(2):154, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900785.

Anonymous:1977:BRG

[Ano77f] Anonymous. Book review: Generation, Compilation, Evaluation & Dissemination of Data for Science & Technology: by B. Dreyfus. Pergamon Press, 1975. 171 pp. $25.00. Computers and Mathematics with Applications, 3 (2):153–154, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122177900773.

Anonymous:1977:BRN

[Ano77g] Anonymous. Book review: Numerical Methods in Finite Element Analysis: by Bathe and Wilson. Prentice–Hall, Englewood Cliffs, New Jersey, June 1976. 521 pp. $28.95. Computers and Mathematics with Applications, 3(2):154, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900797.

Anonymous:1977:BRP

[Ano77h] Anonymous. Book review: Programs and Machines: by Richard Bird. John Wiley, New York, August 1976. 214 pp. $16.95. Computers and Mathematics with Applications, 3(2):153, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900761.

Anonymous:1977:EB

[Ano77i] Anonymous. Editorial board. Computers and Mathematics with Applications, 3(1):ifc, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177901080.

Anonymous:1977:ON

[Ano77j] Anonymous. Obituary notice.

Computers and Mathematics with Applications, 3(2):85, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900591.

Anonymous:1977:OBR

[Ano77k] Anonymous. Other books received. Computers and Mathematics with Applications, 3 (2):155, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900815.

Anonymous:1978:BR

[Ano78a] Anonymous. Books received.

Computers and Mathematics with Applications, 4(3):286, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900482.

Anonymous:1978:EB

[Ano78b] Anonymous. Editorial board.

Computers and Mathematics with Applications, 4(1):ifc, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900019.

Anonymous:1978:E

[Ano78c] Anonymous. Erratum. Computers and Mathematics with Applications, 4(3):ii, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900500.

Anonymous:1978:TIS

[Ano78d] Anonymous. Tenth International Symposium on Mathematical Programming: Montréal 27– 31 August 1979. Computers and Mathematics with Applications, 4(3):i, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900494.

Anonymous:1979:EB

[Ano79a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 5(1):ifc, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901322.

Anonymous:1979:OBR

[Ano79b] Anonymous. Other books received. Computers and Mathematics with Applications, 5 (2):155, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900701.

Anonymous:1979:SIS

[Ano79c] Anonymous. Second International Symposium on Innovative numerical analysis in applied engineering science: Montréal, Canada 16–20 June 1980. Computers and Mathematics with Applications, 5(1):77, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901437.

Arato:1975:DAM

[Ara75] Mátyás Arató. Diffusion approximation for multiprogrammed computer systems. Computers and Mathematics with Applications, 1(3–4):315–326, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900310.

Asai:1976:RRC

[Asa76] H. Asai. A recursive radix conversion formula and its application to multiplication and division. Computers and Mathematics with Applications, 2(3– 4):255–265, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900183.

Balazs:1975:VNF

[Bal75] N. L. Balazs. Viscous nozzle flow as a Sturm–Liouville problem. Computers and Mathematics with Applications, 1(3– 4):291–297, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900280.

Bates:1978:FTS

[Bat78] J. B. Bates. Fourier transform spectroscopy. Computers and Mathematics with Applications, 4(2):73–84, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900202.

Blakley:1979:RSA

[BB79] G. R. Blakley and I. Borosh. Rivest–Shamir–Adleman public key cryptosystems do not always conceal messages. Computers and Mathematics with Applications, 5 (3):169–178, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900397.

Bamberger:1978:PCV

[BJY78] A. Bamberger, J. Jaffre, and J. P. Yvon. Punctual control of a vibrating string: Numerical analysis. Computers and Mathematics with Applications, 4 (2):113–138, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122178900226.

Bellman:1975:DQS

[BKLV75a] R. Bellman, Bayesteh Kashef, E. S. Lee, and R. Vasudevan. Differential quadrature and splines. Computers and Mathematics with Applications, 1(3– 4):371–376, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900383.

Bellman:1975:SHP

[BKLV75b] Richard Bellman, Bayesteh Kashef, E. Stanley Lee, and R. Vasudevan. Solving hard problems by easy methods: Differential and integral quadrature. Computers and Mathematics with Applications, 1 (1):133–143, January 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900139.

Brown:1979:SAN

[Bro79] R. Leonard Brown. Stability analysis of nonlinear differential sequences generated numerically. Computers and Mathematics with Applications, 5(3):187–192, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122179900415.

Butler:1975:LL

[But75] R. Butler. The Lanczos legend. Computers and Mathematics with Applications, 1(??):258–??, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Cooper:1975:NLI

[CC75] Leon Cooper and Mary W. Cooper. Non-linear integer programming. Computers and Mathematics with Applications, 1 (2):215–222, June 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900206.

Cooper:1976:EAC

[CC76] Leon Cooper and Mary W. Cooper. An exact algorithm for the concave transportation problem. Computers and Mathematics with Applications, 2 (1):49–61, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900055.

Cekirge:1978:PFF

[Cek78] H. M. Cekirge. Propagation of fire fronts in forests. Computers and Mathematics with Applications, 4(4):325–332, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900135.

Chaitin:1976:AES

[Cha76] Gregory J. Chaitin. Algorithmic entropy of sets. Computers and Mathematics with Applications, 2(3–4):233–245, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812217690016X.

Choi:1977:BRI

[Cho77a] S. C. Choi. Book review: Introduction to Statistics and Computer Programming: by C. F. Kossack and C. I. Henschke, Holden-Day, San Francisco, 1975. 651 pp. $10.95. Computers and Mathematics with Applications, 3 (2):151–152, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900724.

Choi:1977:BRM

[Cho77b] S. C. Choi. Book review: Multiple Regression and Analysis of Variance: by G. O. Wesolowsky, Wiley Interscience, New York, 1976. 292 pp. $19.95. Computers and Mathematics with Applications, 3(2):150–151, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900700.

Choi:1977:TSS

[Cho77c] S. C. Choi. Two-sample sequential test of normal distributions when one of the sample sizes is fixed. Computers and Mathematics with Applications, 3 (2):125–129, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900657.

Choi:1978:BRMa

[Cho78a] S. C. Choi. Book review: Methods for statistical data analysis of multivariate observations: By G. Gnanadesikan, Wiley Interscience, New York, 1977, 311 pp. Computers and Mathematics with Applications, 4(2):199, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900299.

Choi:1978:BRMb

[Cho78b] S. C. Choi. Book review: Methods for Statistical Data Analysis of Multivariate Observations: By G. Gnanadesikan. Wiley Interscience New York, 1977, 311 pp. Computers and Mathematics with Applications, 4(4):378, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900196.

Cocke:1976:PCP

[CK76] John Cocke and Ken Kennedy. Profitability computations on program flow graphs. Computers and Mathematics with Applications, 2(2):145–159, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900274.

Claessens:1976:RHI

[Cla76] G. Claessens. The rational Hermite interpolation problem and some related recurrence formulas. Computers and Mathematics with Applications, 2(2):117–123, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122176900237.

Chhikara:1975:DSM

[CO75] Raj S. Chhikara and Patrick L.

Odell. On designing simulation models for evaluating discriminant analysis routines. Computers and Mathematics with Applications, 1(1):69–78, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900085.

Cooper:1975:FCP

[Coo75] Leon Cooper. The fixed charge problem — I: a new heuristic method. Computers and Mathematics with Applications, 1 (1):89–95, January 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900103.

Cooper:1977:DOF

[Coo77a] Leon Cooper. The determination of optimal faculty staffing for a university. Computers and Mathematics with Applications, 3 (3):161–174, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900918.

Cooper:1977:BRQ

[Coo77b] Robert B. Cooper. Book review: Queueing Systems, Volume II: Computer Applications: By Leonard Kleinrock. Wiley– Interscience, New York, 1976, xx + 549 pp. Computers and Mathematics with Applications, 3(1):83, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177901183.

Cooper:1978:BRC

[Coo78a] Leon Cooper. Book review: Combinatorial algorithms: Theory and practice: By E. M. Reingold, J. Nievergelt and N. Deo. Prentice–Hall, Englewood Cliffs, NJ, 1977, $18.95. Computers and Mathematics with Applications, 4(3):283, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900433.

Cooper:1978:BRS

[Coo78b] Leon Cooper. Book review: Software metrics: By Tom Gilb. Computers and Mathematics with Applications, 4(3):283, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900421.

Cooper:1978:STL

[Coo78c] Leon Cooper. The stochastic transportation-location problem. Computers and Mathematics with Applications, 4(3):265–275, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812217890038X.

Cooper:1979:IAN

[Coo79] Mary W. Cooper. An improved algorithm for non-linear integer programming. Computers and Mathematics with Applications, 5 (2):145–152, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900671.

Cooper:1977:FHS

[CS77] Leon Cooper and Bruce W.

Schmeiser. A family of heuristic solution methods for a class of nonlinear programming problems. Computers and Mathematics with Applications, 3(1):23–39, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122177901110.

Cekirge:1978:CCF

[CU78] H. M. Cekirge and E. Ural. Critical coolant flow velocities in reactors having parallel fuel plates. Computers and Mathematics with Applications, 4(2):153–156, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122178900251.

Delforge:1978:ACC

[Del78] Jacques Delforge. Application of the concept of the conditioning of a matrix to an identification problem. Computers and Mathematics with Applications, 4 (2):139–148, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900238.

Davis:1975:ESL

[DGN75] W. R. Davis, L. H. Green, and L. K. Norris. Elucidation of some of Lanczos’ thoughts concerning spacetime curvature and the local and non-local parts of the Riemann tensor. Computers and Mathematics with Applications, 1 (3–4):361–364, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/089812217590036X.

Dimsdale:1977:BPB

[Dim77] B. Dimsdale. Bicubic patch bounds. Computers and Mathematics with Applications, 3(2): 95–104, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812217790061X.

Dinkel:1976:MLC

[DK76] John J. Dinkel and Gary A.

Kochenberger. Maximization of linearly constrained polynomials. Computers and Mathematics with Applications, 2(2):101–110, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122176900213.

Decell:1977:FCD

[DM77] Henry P. Decell, Jr. and Shailesh M.

Mayekar. Feature combinations and the divergence criterion. Computers and Mathematics with Applications, 3(1):71–76, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812217790116X.

Dang:1978:FEM

[DN78] Dac Q. Dang and D. H. Norrie. A finite element method for the solution of singular integral equations. Computers and Mathematics with Applications, 4 (3):219–224, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900330.

Davis:1979:MET

[DS79] Martin Davis and Jacob T. Schwartz. Metamathematical extensibility for theorem verifiers and proof-checkers. Computers and Mathematics with Applications, 5(3):217–230, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900440.

deVeubeke:1975:NIL

[dVDB75] B. Fraeijs de Veubeke and C. Delcourt-Bon. The numerical integration of laminar boundary layer equations. Computers and Mathematics with Applications, 1(2):167–173, June 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900164.

Dymond:1979:IAT

[Dym79] Patrick W. Dymond. Indirect addressing and the time relationships of some models of sequential computation. Computers and Mathematics with Applications, 5 (3):193–209, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900427.

Erdos:1975:SGD

[EGS75] P. Erdős, R. L. Graham, and E. Szemerédi. On sparse graphs with dense long paths. Computers and Mathematics with Applications, 1(3–4):365–369, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900371.

Evans:1978:PCE

[EH78] D. J. Evans and M. Hatzopoulos. The parallel calculation of the eigenvalues of a real matrix A. Computers and Mathematics with Applications, 4(3): 211–218, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900329.

Elliott:1976:BRI

[Ell76] D. L. Elliott. Book review: Introduction to Mathematical Biology: By S. I. Rubinow. Wiley– Interscience, New York, 1975, 386 pp. $22. Computers and Mathematics with Applications, 2(3–4):267, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900195.

Evans:1979:GSF

[EO79] D. J. Evans and S. O. Okolie. A generalised sparse factorisation method for the solution of periodic tridiagonal systems. Computers and Mathematics with Applications, 5(3):211–216, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122179900439.

Erdos:1979:EC

[ES79] Paul Erdős and Joel Spencer. Evolution of the n-cube. Computers and Mathematics with Applications, 5(1):33–39, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181901371.

Farlow:1977:ASS

[Far77] S. J. Farlow. Analytical solutions to second order differential equations. Computers and Mathematics with Applications, 3(1):15–22, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177901109.

Fatunla:1976:NAN

[Fat76] Simeon Ola Fatunla. A new algorithm for numerical solution of ordinary differential equations. Computers and Mathematics with Applications, 2(3– 4):247–253, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900171.

Fatunla:1978:VOO

[Fat78] Simeon O. Fatunla. A variable order one-step scheme for numerical solution of ordinary differential equations. Computers and Mathematics with Applications, 4(1):33–41, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900056.

Fair:1978:CFS

[FD78] Wyman Fair and Daniel A. Doyle.

Continued fraction solution to a nonlinear integral equation. Computers and Mathematics with Applications, 4(1):11–14, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900032.

Franklin:1976:SDD

[FG76] M. A. Franklin and R. K. Gupta.

Stack distance density and page fault probability in virtual memory systems. Computers and Mathematics with Applications, 2(1):41–48, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122176900043.

Fix:1977:DBC

[FG77] George Fix and Max Gunzburger. Downstream boundary conditions for viscous flow problems. Computers and Mathematics with Applications, 3(1):53–63, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177901146.

Fix:1979:FEM

[FGN79] George J. Fix, Max D. Gunzburger, and R. A. Nicolaides. On finite element methods of the least squares type. Computers and Mathematics with Applications, 5(2):87–98, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122179900622.

Fix:1977:FEA

[FN77] G. J. Fix and B. Neta. Finite element approximation of a nonlinear diffusion problem. Computers and Mathematics with Applications, 3(4):287–298, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900852.

Fraenkel:1977:PAG

[Fra77] Aviezri S. Fraenkel. The particles and antiparticles game. Computers and Mathematics with Applications, 3(4):327–328, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900906.

Fraenkel:1975:SCG

[FT75] A. S. Fraenkel and U. Tassa.

Strategy for a class of games with dynamic ties. Computers and Mathematics with Applications, 1(2):237–254, June 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812217590022X.

Gargantini:1976:CPN

[Gar76] Irene Gargantini. Comparing parallel Newton’s method with parallel Laguerre’s method. Computers and Mathematics with Applications, 2(3–4):201–206, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900122.

Gargantini:1979:NSS

[Gar79] Irene Gargantini. The numerical stability of simultaneous iterations via square-rooting. Computers and Mathematics with Applications, 5(1):25–31, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812218190136X.

Gerstring:1977:NME

[Ger77] John M. Gerstring, Jr. Numerical methods for eigensystems: the Orr–Sommerfeld problem. Computers and Mathematics with Applications, 3(1):47–52, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177901134.

Gehrlein:1979:PPM

[GF79] William V. Gehrlein and Peter C. Fishburn. Proportions of profiles with a majority candidate. Computers and Mathematics with Applications, 5(2):117–124, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122179900646.

Georg:1976:SSA

[GK76] D. D. Georg and R. F. Keller. Subspace selection algorithms to be used with the nonlinear projection methods in solving systems of nonlinear equations. Computers and Mathematics with Applications, 2(3–4):225–232, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900158.

Gabay:1976:DAS

[GM76] Daniel Gabay and Bertrand Mercier. A dual algorithm for the solution of nonlinear variational problems via finite element approximation. Computers and Mathematics with Applications, 2(1):17–40, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900031.

Gout:1979:CHR

[Gou79a] J. L. Gout. Construction of a Hermite rational “Wachspress type” finite element. Computers and Mathematics with Applications, 5 (4):337–347, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900932.

Gout:1979:IEE

[Gou79b] J. L. Gout. Interpolation error estimates on Hermite rational “Wachspress type” third degree finite element. Computers and Mathematics with Applications, 5(4):349–357, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900944.

George:1978:ADA

[GPV78] Alan George, W. G. Poole, Jr., and R. G. Voigt. Analysis of dissection algorithms for vector computers. Computers and Mathematics with Applications, 4 (4):287–304, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812217890010X.

Greenspan:1977:AAM

[Gre77] Donald Greenspan. Arithmetic applied mathematics. Computers and Mathematics with Applications, 3(4):253–270, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900839.

Geisler:1975:PEB

[GTG75] E. G. Geisler, A. A. Tal, and D. P. Garg. On the a-posteriori error bounds for the solution of ordinary nonlinear differential equations. Computers and Mathematics with Applications, 1(3– 4):407–416, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900425.

Haimo:1978:RLT

[Hai78] Deborah Tepper Haimo. Representations for Laguerre temperatures. Computers and Mathematics with Applications, 4(2): 173–187, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900275.

Horvay:1977:CLM

[Hor77] Gabriel Horvay. Cornelius Lanczos, in memoriam. Computers and Mathematics with Applications, 3(4):251–252, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900827.

Hallum:1975:CAM

[HP75] C. R. Hallum and M. D. Pore.

Computational aspects of matrix generalized inversion for the computer with applications. Computers and Mathematics with Applications, 1(1):145–150, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122175900140.

Hill:1977:CSI

[HS77] H. W. Hill and W. Squire. Computation of semi-infinite integrals on a programmable hand calculator. Computers and Mathematics with Applications, 3(3): 191–196, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900943.

Judelman:1977:EDG

[JJ77] M. Judelman and D. H. Jacobson. The effect of data grid size on certain interpolation methods for unconstrained function minimization. Computers and Mathematics with Applications, 3 (3):175–182, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812217790092X.

Jain:1979:CST

[JL79] P. C. Jain and B. L. Lohar. Cubic spline technique for coupled non-linear parabolic equations. Computers and Mathematics with Applications, 5(3): 179–185, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900403.

Jambekar:1978:CEA

[JS78a] Anil B. Jambekar and David I. Steinberg. Computational experience with an algorithm for 0– 1 integer programming. Computers and Mathematics with Applications, 4(4):305–314, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900111.

Jambekar:1978:IEA

[JS78b] Anil B. Jambekar and David I. Steinberg. An implicit enumeration algorithm for the all integer programming problem. Computers and Mathematics with Applications, 4(1):15–31, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900044.

Keats:1975:HML

[Kea75] R. G. Keats. Herbert Melvin Lieberstein in memoriam. Computers and Mathematics with Applications, 1(1):41–42, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122175900061.

Katz:1977:OSP

[KFS77] I. Norman Katz, Mark A.

Franklin, and A. Sen. Optimally stable parallel predictors for Adams–Moulton correctors. Computers and Mathematics with Applications, 3(3):217–233, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122177900979.

Kintigh:1977:BRM

[Kin77] Keith Kintigh. Book review:

Mathematics and Computers in Archaeology: By J. E. Doran and F. R. Hodson. Harvard University Press, Cambridge, Massachusetts. 1975. 381 pp. $18.00. Computers and Mathematics with Applications, 3(2):149–150, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900694.

Kalaba:1978:IMM

[KM78] R. Kalaba and K. Mease. An imbedding method for matrix eigenvalue problems. Computers and Mathematics with Applications, 4(1):53–59, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812217890007X.

Kelman:1977:ACD

[KMS77] Robert B. Kelman, Johnathan P. Maheffy, and J. Timothy Simpson. Algorithms for classic dual trigonometric equations. Computers and Mathematics with Applications, 3(3):203–215, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900967.

Knuth:1975:SCS

[Knu75] Elöd Knuth. A structured computer system model. Computers and Mathematics with Applications, 1(3–4):327–336, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900322.

Knuth:1976:EPC

[Knu76] Donald E. Knuth. Evaluation of Porter’s constant. Computers and Mathematics with Applications, 2(2):137–139, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900250.

Kotiah:1979:BRI

[Kot79] Thoddi C. T. Kotiah. Book review: Introductory Applied Statistics in Science: by Sung C. Choi, Prentice–Hall, Englewood Cliffs, New Jersey, 1978. 278 pp. $15.95. Computers and Mathematics with Applications, 5(2):153–154, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122179900695.

Katz:1978:NVC

[KPR78] I. Norman Katz, Alberto G.

Peano, and Mark P. Rossow. Nodal variables for complete conforming finite elements of arbitrary polynomial order. Computers and Mathematics with Applications, 4(2):85–112, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900214.

Kelman:1975:ASD

[KS75a] Robert B. Kelman and J. Timothy Simpson. An algorithm for solving a dual cosine series. Computers and Mathematics with Applications, 1(2):195–199, June 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122175900188.

Kennedy:1975:IST

[KS75b] K. Kennedy and J. Schwartz. An introduction to the set theoretical language SETL. Computers and Mathematics with Applications, 1(1):97–119, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900115.

Kalaba:1978:NAE

[KS78a] R. E. Kalaba and K. Spingarn. Numerical approaches to the eigenvalues of Saaty’s matrices for fuzzy sets. Computers and Mathematics with Applications, 4(4):369–375, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900172.

Kalaba:1978:PBB

[KS78b] R. E. Kalaba and K. Spingarn. Post buckling beam configurations via an imbedding method. Computers and Mathematics with Applications, 4(1):1–10, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122178900020.

Kotiah:1978:BRS

[KS78c] Thoddi C. T. Kotiah and David I. Steinberg. Book review: Statistical methods for digital computers: Edited by Kurt Enslein, Anthony Ralston and Herbert S. Wilf, Wiley Interscience, New York, 1977, 454 pp. $24.95. Computers and Mathematics with Applications, 4 (2):199–200, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900305.

Kalaba:1975:IMB

[KSZ75] R. E. Kalaba, K. Spingarn, and E. Zagustin. An imbedding method for buckling loads. Computers and Mathematics with Applications, 1(3–4):277–284, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122175900267.

Kalaba:1975:PRA

[KZ75] R. E. Kalaba and E. A. Zagustin. On penetration of a rigid axisymmetric punch into an elastic layer. Computers and Mathematics with Applications, 1 (1):13–25, January 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900048.

Kalaba:1978:EPB

[KZ78] R. Kalaba and E. Zagustin. Estimation of parameters for a beam on an elastic foundation with large deflections and built in at both ends. Computers and Mathematics with Applications, 4 (3):257–263, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900378.

Kalaba:1977:MDM

[KZHH77] Robert E. Kalaba, Elena Zagustin, W. Holbrow, and R. Huss. A modification of Davidenko’s method for nonlinear systems. Computers and Mathematics with Applications, 3(4):315–319, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122177900888.

Kalaba:1978:IVM

[KZM78] Robert E. Kalaba, Elena A. Zagustin, and George Meier. An initial value method for orthogonal polynomials. Computers and Mathematics with Applications, 4 (2):189–197, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900287.

LeBlanc:1979:EDA

[LA79] Larry J. LeBlanc and Mustafa Abdulaal. An efficient dual approach to the urban road network design problem. Computers and Mathematics with Applications, 5(1):11–19, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901346.

LeBlanc:1977:HAL

[LeB77] Larry J. LeBlanc. A heuristic approach for large scale discrete stochastic transportationlocation problems. Computers and Mathematics with Applications, 3(2):87–94, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900608.

Luke:1975:PCF

[LFW75] Yudell L. Luke, Wyman Fair, and Jet Wimp. Predictor– corrector formulas based on rational interpolants. Computers and Mathematics with Applications, 1(1):3–12, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900036.

Lanczos:1975:FAR

[LG75] Cornelius Lanczos and Barbara Gellai. Fourier analysis of random sequences. Computers and Mathematics with Applications, 1 (3–4):269–276, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900255.

Lieberstein:1975:HOA

[LI75] H. M. Lieberstein and C. D. Isaacs. Hybrid optimization of accelerated successive replacements for linear and nonlinear systems. Computers and Mathematics with Applications, 1(1):27–40, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812217590005X.

Lieberstein:1975:MTS

[Lie75] H. Melvin Lieberstein. Mathematical treatment of some problems in physiology. Computers and Mathematics with Applications, 1(1):43–67, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900073.

Lee:1976:NBV

[LL76] E. S. Lee and P. Lee. Nonlinear boundary value problems and multicomponent distillation. Computers and Mathematics with Applications, 2(1):85–93, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122176900080.

Lee:1975:OWF

[LM75] E. Stanley Lee and P. K. Misra. Optimal weighting function for the invariant imbedding estimator. Computers and Mathematics with Applications, 1(1):

79–88, January 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900097.

Lee:1978:CNM

[LP78] Ding Lee and Stanley Preiser.

A class of nonlinear multistep A-stable numerical methods for solving stiff differential equations. Computers and Mathematics with Applications, 4(1):43–51, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122178900068.

Lii:1975:ABS

[LR75] Keh-Shin Lii and M. Rosenblatt.

Asymptotic behavior of a spline estimate of a density function. Computers and Mathematics with Applications, 1(2):223–235, June 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122175900218.

Luke:1975:SRU

[Luk75] Y. L. Luke. Some remarks on uniform asymptotic expansions for Bessel functions. Computers and Mathematics with Applications, 1 (3–4):285–290, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900279.

Luke:1977:EPA

[Luk77] Yudell L. Luke. On the error in Padé approximations for functions defined by Stieltjes integrals. Computers and Mathematics with Applications, 3(4): 307–314, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900876.

Lee:1976:IIM

[LVW76] E. S. Lee, R. Vasudevan, and K. M. Wang. Invariant imbedding, method of characteristics, and parameter estimation. Computers and Mathematics with Applications, 2(1):95–98, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900092.

Marsaglia:1977:SMG

[Mar77] George Marsaglia. The squeeze method for generating gamma variates. Computers and Mathematics with Applications, 3(4): 321–325, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812217790089X.

McConnell:1975:CLD

[McC75] J. R. McConnell. Cornelius Lanczos in Dublin (1953–1974). Computers and Mathematics with Applications, 1(??):263–264, ????

1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

McLeod:1979:EGR

[McL79a] Robin J. Y. McLeod. Extensions and generalisations of rational and high order transformation methods. Computers and Mathematics with Applications, 5 (4):267–276, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900877.

McLeod:1979:F

[McL79b] Robin J. Y. McLeod. Foreword.

Computers and Mathematics with Applications, 5(4):iii–iv, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/089812217990083X.

McLeod:1979:STD

[McL79c] Robin J. Y. McLeod. On the stability of two-dimensional interpolation and high order bases for curved finite elements. Computers and Mathematics with Applications, 5(4):249–265, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900865.

Miller:1975:TPW

[Mil75] John J. H. Miller. On the type of a polynomial f, when f^{∗}(0)f − f(0)f^{∗} is a monomial. Computers and Mathematics with Applications, 1(3–4):299–306, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900292.

Mitchell:1979:ACA

[Mit79]  A. R. Mitchell.  Advantages

of cubics for approximating element boundaries. Computers and Mathematics with Applications, 5(4):321–327, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900919.

McLeod:1979:PPA

[MM79a] Robin J. Y. McLeod and A. R. Mitchell. A piecewise parabolic C^{1} approximation technique for curved boundaries. Computers and Mathematics with Applications, 5(4):277–284, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900889.

McLeod:1979:NCH

[MM79b] Robin J. Y. McLeod and R. Murphy. A numerical comparison of high order transformation and isoparametric transformation methods. Computers and Mathematics with Applications, 5 (4):241–247, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900853.

Miranker:1976:EGS

[MP76] W. L. Miranker and O. Pironneau. An example of a global shock fitting method. Computers and Mathematics with Applications, 2(1):63–71, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900067.

Mathis:1976:NSH

[MR76] F. H. Mathis and G. W. Reddien. Numerical solution of hereditary control problems using necessary conditions. Computers and Mathematics with Applications, 2 (3–4):193–200, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900110.

Mullineux:1977:NII

[MR77] N. Mullineux and J. R. Reed. Numerical inversion of integral transforms. Computers and Mathematics with Applications, 3(4): 299–306, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900864.

Metropolis:1977:SAP

[MT77] N. Metropolis and Stephen M. Tanny. Significance arithmetic: the probability of carrying. Computers and Mathematics with Applications, 3(1):77–81, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177901171.

Miele:1975:HAO

[MTC75] A. Miele, J. L. Tietze, and J. R. Cloutier. A hybrid approach to optimal control problems with bounded state. Computers and Mathematics with Applications, 1(2):175–194, June 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900176.

Nahin:1976:EAA

[Nah76] Paul J. Nahin. An error analysis of an algorithm for Marcum’s Qfunction. Computers and Mathematics with Applications, 2(3– 4):207–210, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900134.

Neta:1978:FEA

[Net78] B. Neta. Finite element approximation of a nonlinear parabolic problem. Computers and Mathematics with Applications, 4(3):

247–255, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900366.

Neuts:1975:CUM

[Neu75] Marcel F. Neuts. Computational uses of the method of phases in the theory of queues. Computers and Mathematics with Applications, 1(2):151–166, June 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900152.

Ng:1977:CAL

[Ng77] Edward W. Ng. Computations and applications of linear hypergeometric transformations. Computers and Mathematics with Applications, 3(1):65–70, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177901158.

Nievergelt:1975:CME

[Nie75] Jurg Nievergelt. Computers and mathematics education. Computers and Mathematics with Applications, 1(1):121–132, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122175900127.

Overbeek:1976:CRE

[OMW76] R. Overbeek, J. McCharen, and L. Wos. Complexity and related enhancements for automated theorem-proving programs. Computers and Mathematics with Applications, 2(1):1–16, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812217690002X.

Ornstein:1977:BRS

[Orn77] W. Ornstein. Book review: Some modern mathematics for physicists and other outsiders: By Paul Roman. Pergamon Press, New York, 1975. Volume 1: 378 pp., $25.00; Volume 2: 335 pp., $25.00. Computers and Mathematics with Applications, 3(1):83–84, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177901195.

Orr:1979:BRD

[Orr79a] Patrick K. Orr. Book review: Discrete Mathematics in Computer Science: By Donald F. Stanat and David F. McAllister. Prentice–Hall, Englewood Cliffs, N.J. 1977, 401 pp., $17.50. Computers and Mathematics with Applications, 5(2):153, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900683.

Orr:1979:BRC

[Orr79b] Patrick K. Orr. Book review:

The complexity of Computing: By John E. Savage. Wiley, New York, 1976. $22.95. Computers and Mathematics with Applications, 5(1):76, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901425.

Ortiz:1975:SST

[Ort75a] E. L. Ortiz. Step by step Tau method. Part I. Piecewise polynomial approximations. Computers and Mathematics with Applications, 1(3–4):381–392, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900401.

Ortiz:1975:LIA

[Ort75b] Eduardo L. Ortiz. Lanczos and the Institute for Advanced Studies in the early sixties. Computers and Mathematics with Applications, 1(??):266–??, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Ostrowski:1975:EAF

[Ost75] A. M. Ostrowski. An estimate of the approximation to a fixed point of an operator. Computers and Mathematics with Applications, 1(3–4):427, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900449.

Overbeek:1975:IHR

[Ove75] Ross A. Overbeek. An implementation of hyper-resolution. Computers and Mathematics with Applications, 1(2):201–214, June 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/089812217590019X.

Peano:1976:HCF

[Pea76] Alberto Peano. Hierarchies of conforming finite elements for plane elasticity and plate bending. Computers and Mathematics with Applications, 2(3– 4):211–224, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900146.

Petruska:1978:FEC

[Pet78] G. Petruska. Finite element convergence on a fixed grid. Computers and Mathematics with Applications, 4(1):67–71, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900093.

Pollack:1976:BRC

[Pol76] Seymour V. Pollack. Book review: Computers Communications and Society: by Murray Laver. Oxford University Press, 1975. 99pp. $11.25. Computers and Mathematics with Applications, 2(2):190, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900316.

Rao:1978:EFC

[Rao78] T. Mahadeva Rao. Error-free computation of characteristic polynomial of a matrix. Computers and Mathematics with Applications, 4(1):61–65, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900081.

Rodin:1975:MCL

[RBE^{+}75] Ervin Y. Rodin, R. Butler, A. Erdélyi, B. Gellai, J. R. Mc-Connell, Eduardo L. Ortiz, and I. Rhodes. In memory of Cornelius Lánczos. Computers and Mathematics with Applications, 1 (3–4):257–268, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900243.

Robers:1976:SFC

[RC76] Philip Robers and Leon Cooper.

A study of the fixed charge transportation problem. Computers and Mathematics with Applications, 2(2):125–135, ???? 1976.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900249.

Rogers:1977:NLD

[RC77] C. Rogers and H. M. Cekirge. Non-linear dielectrics. Computers and Mathematics with Applications, 3(2):105–112, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900621.

Reddy:1977:EUS

[Red77] J. N. Reddy. Existence and uniqueness of solutions to a stationary finite element model of the biharmonic equation. Computers and Mathematics with Applications, 3(2):135–147, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122177900670.

Reddien:1979:NMH

[Red79] G. W. Reddien. Newton’s method and high order singularities. Computers and Mathematics with Applications, 5(2):79–86, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122179900610.

Rhodes:1977:CDH

[Rho77] Ida Rhodes. Computation of the dates of the Hebrew New Year and Passover. Computers and Mathematics with Applications, 3 (3):183–190, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900931.

Ries:1976:APS

[Rie76] John Ries. An algorithm for the power series solution to the first order differential equation where the right side is the ratio of two polynomials. Computers and Mathematics with Applications, 2(2):141–143, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122176900262.

Reddy:1977:NOE

[RO77] C. T. Reddy and J. T. Oden.

A note on optimal error estimates for finite element approximations of a nonlinear twopoint boundary-value problem. Computers and Mathematics with Applications, 3(2):117–123, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122177900645.

Rodin:1975:F

[Rod75] Ervin Y. Rodin. Foreword. Computers and Mathematics with Applications, 1(1):1, January 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900024.

Rodin:1977:BRN

[Rod77a] Ervin Y. Rodin. Book review: Nonlinear Programming Analysis and Methods: by Mordecal Avriel. Prentice–Hall, New York (July 1976) 512 pp. $24.95. Computers and Mathematics with Applications, 3(2):151, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900712.

Rodin:1977:BRH

[Rod77b] Ervin Y. Rodin. Book review: The Heat Equation: by D. V. Widder, Academic Press, New York, 1975. 267 pp. $22.50. Computers and Mathematics with Applications, 3(2):149, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900682.

Rosser:1975:NPD

[Ros75] J. Barkley Rosser. Nine-point difference solutions for Poisson’s equation. Computers and Mathematics with Applications, 1(3– 4):351–360, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122175900358.

Rosen:1978:OMP

[Ros78] Bernard Rosen. Operator method for the perturbative solution of ordinary differential equations. Computers and Mathematics with Applications, 4(3):225–235, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122178900342.

Rodabaugh:1977:CMI

[RT77] David Rodabaugh and Skip Thompson. Corrector methods with increased ranges of stability. Computers and Mathematics with Applications, 3(3):197–201, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122177900955.

Rodabaugh:1978:ATM

[RT78] D. J. Rodabaugh and Skip Thompson. Adams-type methods with increased ranges of stability. Computers and Mathematics with Applications, 4(4): 349–357, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900159.

Ruas:1979:AGT

[Rua79] Vitoriano Ruas. Automatic generation of triangular finite element meshes. Computers and Mathematics with Applications, 5 (2):125–140, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900658.

Sachdev:1976:CES

[Sac76] P. L. Sachdev. A class of exact solutions of boundary value problems for Burger’s equation. Computers and Mathematics with Applications, 2(2):111–116, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122176900225.

Shampine:1977:DSF

[SH77] L. F. Shampine and K. L. Hiebert. Detecting stiffness with the Fehlberg (4, 5) formulas. Computers and Mathematics with Applications, 3(1):41–46, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177901122.

Shampine:1979:BSO

[Sha79] Lawrence F. Shampine. Better software for ODEs. Computers and Mathematics with Applications, 5(3):157–161, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900373.

Sheu:1978:TAM

[She78] Ming-Goei Sheu. On theories and applications of mixed finite element methods for linear boundary-value problems. Computers and Mathematics with Applications, 4(4):333–347, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122178900147.

Silliman:1979:CSQ

[Sil79]  Sherwood D. Silliman.  On

complete semicardinal quadrature formulae. Computers and Mathematics with Applications, 5(1):41–49, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901383.

Sims:1978:NSC

[Sim78] S. E. Sims. A note on some convergence properties of spline functions. Computers and Mathematics with Applications, 4(3): 277–279, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900391.

Smith:1978:OPM

[Smi78] A. C. Smith. A one-parameter method for accelerating the convergence of sequences and series. Computers and Mathematics with Applications, 4(2):157–171, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122178900263.

Stephanou:1976:BRL

[SS76] Harry E. Stephanou and George N. Saridis. Book review: Legged Locomotion Robots and Anthropomorphic Mechanisms: by M. Vukobratovic. Mihailo Pupin Institute, Belgrade, 1975, 346 pp. $23. Computers and Mathematics with Applications, 2(2):189, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122176900304.

Slotnick:1978:NCC

[SS78] D. L. Slotnick and Ahmed Sameh. Numerical calculation and computer design. Computers and Mathematics with Applications, 4 (3):201–210, ???? 1978. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900317.

Steinberg:1978:BRL

[Ste78a] D. I. Steinberg. Book review: Linear and combinatorial programming: By Katta G. Murty. John Wiley, New York, 1976, 567 pp. Computers and Mathematics with Applications, 4(3):281, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900408.

Steinberg:1978:BRMb

[Ste78b] D. I. Steinberg. Book review:

Mathematical models for surface water hydrology: Edited by T. A. Ciriani, U. Maione and J. R. Wallis. Wiley Interscience, New York, 1977, 423 pp., $27.50. Computers and Mathematics with Applications, 4(3):285, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900469.

Steinberg:1978:BRO

[Ste78c] D. I. Steinberg. Book review: Operations research handbook: Standard algorithms and methods: By Horst A. Eiselt and Helmut von Frajer. De Gruyter. Berlin, 1977, 398 pp., $32.90. Computers and Mathematics with Applications, 4(3):282, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812217890041X.

Steinberg:1978:BRC

[Ste78d] D. I. Steinberg. Book review: The complexity of computing: By John E. Savage. Wiley Interscience, New York, 1976 391 pp., $22.95. Computers and Mathematics with Applications, 4(3):285, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900470.

Steinberg:1978:BRMa

[Ste78e] D. I. Steinberg. Book review: The mini-computer in the laboratory: With examples using the PDP-11: By James W. Cooper. Wiley Interscience, New York, 1977, 365 pp., $19.50. Computers and Mathematics with Applications, 4(3):285, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900457.

Stoutemyer:1979:SCV

[Sto79] David R. Stoutemyer. Symbolic computer vector analysis. Computers and Mathematics with Applications, 5(1):1–9, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181901334.

Strand:1976:SAB

[Str76] O. N. Strand. Some aspects of the behavior of regularized solutions as the amount of smoothing is varied. Computers and Mathematics with Applications, 2 (2):181–187, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900298.

Szabo:1979:SRD

[Sza79] B. A. Szabo. Some recent developments in finite element analysis. Computers and Mathematics with Applications, 5(2):99–115, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122179900634.

Szekeres:1975:CSC

[Sze75] G. Szekeres. Cosmology with spinor connection. Computers and Mathematics with Applications, 1(3–4):345–350, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900346.

Tally:1978:COL

[Tal78] William Tally. On the convergence of optimal linear combination procedures. Computers and Mathematics with Applications, 4(2):149–152, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/089812217890024X.

Tally:1979:CCO

[Tal79] William Tally. A convergence criterion for optimal linear combination procedures. Computers and Mathematics with Applications, 5(1):21–24, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122181901358.

Taub:1975:LSR

[Tau75] A. H. Taub. Lanczos’ splitting of the Riemann tensor. Computers and Mathematics with Applications, 1(3–4):377–380, ???? 1975. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900395.

Thomas:1978:CEE

[Tho78] G. Thomas. The consistence error of extended variational principles in the finite element method. Computers and Mathematics with Applications, 4(4):315–324, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

                                        sciencedirect.com/science/
                                        article/pii/0898122178900123.

Tomko:1975:PUS

[Tom75] József Tomkó. Processor utilization study. Computers and Mathematics with Applications, 1(3– 4):337–344, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900334.

Turan:1975:LRM

[Tur75] Paul Turán. On the latent roots of λ matrices. Computers and Mathematics with Applications, 1 (3–4):307–313, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122175900309.

Wachspress:1977:SNS

[Wac77] Eugene L. Wachspress. Stability of numerical solution of semi-discrete parabolic equations. Computers and Mathematics with Applications, 3(2):131–133, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http://www.

sciencedirect.com/science/
article/pii/0898122177900669.

Wachspress:1979:RBF

[Wac79] Eugene L. Wachspress. Rational basis functions for curved elements. Computers and Mathematics with Applications, 5(4): 235–239, ???? 1979. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900841.

Wainwright:1979:ATC

[Wai79] Roger L. Wainwright. Acceleration techniques for a class of x-projection methods for solving systems of linear equations. Computers and Mathematics with Applications, 5(1):59–73, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901401.

Wallace:1978:BRS

[Wal78] Norval D. Wallace. Book review: Simulation with GPSS and GPSS V: By P. A. Bobiller, B. C. Kahan and A. R. Probst. Prentice–Hall, Englewood Cliffs NJ, 1976, 495 pp., $18.50. Computers and Mathematics with Applications, 4(4):377, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122178900184.

Warzee:1975:FEA

[War75] G. Warzee. Finite element analysis of nonlinear magnetic fields. Computers and Mathematics with Applications, 1(3– 4):417–425, ???? 1975. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122175900437.

Warren:1976:DCV

[War76] H. S. Warren, Jr. Definition of the concept “vector” in set theoretic programming languages. Computers and Mathematics with Applications, 2(1):73–83, ???? 1976. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900079.

Weiss:1977:EAC

[Wei77] Joel R. Weiss. An efficient algorithm for capacity state calculations for electric generating systems. Computers and Mathematics with Applications, 3(2): 113–116, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177900633.

Wides:1978:BRS

[Wid78] Jeffrey W. Wides. Book review: Survey and opinion research: Procedures for processing and analysis: By John A. Sonquist and William C. Dunkelberg. Prentice–Hall, Englewood Cliffs, NJ, 1977, 502 pp. Computers and Mathematics with Applications, 4(3):284, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900445.

Wides:1979:BRA

[Wid79] Jeffrey W. Wides. Book review: The Analysis of Cross-Classifications: By H. T. Reynolds. The Free Press, New York, 1977, xiv + 236pp. Computers and Mathematics with Applications, 5(3):232, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900488.

Winker:1976:CDA

[Win76] Steven K. Winker. Complete demodulation for automatic theorem proving. Computers and Mathematics with Applications, 2 (2):161–179, ???? 1976. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122176900286.

Windham:1979:RLD

[Win79] Michael P. Windham. Reliability for linear differential equations with noisy coefficients. Computers and Mathematics with Applications, 5(2):141–144, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/089812217990066X.

Wainwright:1977:APM

[WK77] Roger L. Wainwright and R. F. Keller. Algorithms for projection methods for solving linear systems of equations. Computers and Mathematics with Applications, 3(3):235–245, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.
                                        sciencedirect.com/science/
                                        article/pii/0898122177900980.

Wong:1977:AEH

[Won77] R. Wong. Asymptotic expansions of Hankel transforms of functions with logarithmic singularities. Computers and Mathematics with Applications, 3(4): 271–286, ???? 1977. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

                                        tronic).  URL http://www.

sciencedirect.com/science/
article/pii/0898122177900840.

Werner:1978:NQR

[WW78] H. Werner and L. Wuytack. Nonlinear quadrature rules in the presence of a singularity. Computers and Mathematics with Applications, 4(3):237–245, ???? 1978. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122178900354.

Zeigler:1977:FCB

[ZE77] Bernard P. Zeigler and Daniel Ewenczyk. Formalization and confirmation of the Boyd–Epley operating system model. Computers and Mathematics with Applications, 3(1):1–13, ???? 1977. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122177901092.

Zeigler:1979:BRA

[Zei79a] Bernard P. Zeigler. Book review: Application-Oriented Algebra: By James L. Fisher. IEP, Dun–Donelley, New York 1977. Computers and Mathematics with Applications, 5(3):231, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900464.

Zeigler:1979:BRD

[Zei79b] Bernard P. Zeigler. Book review: Discrete Mathematics in Computer Science: By Donald F. Stanat and David F. McAllister. Prentice Hall, Englewood Cliffs, New Jersey 1977. Computers and Mathematics with Applications, 5(3):231–232, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900476.

Zeigler:1979:BRM

[Zei79c] Bernard P. Zeigler. Book review: Mathematical Modeling: By J. G. Andrews and R. R. McLone. Butterworth, London. Computers and Mathematics with Applications, 5(3):231, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122179900452.

Zeigler:1979:BRF

[Zei79d] Bernard P. Zeigler. Book review: The foundations of cybernetics: By F. H. George. Gordon and Beach, London, 1977. Computers and Mathematics with Applications, 5(1):75, ???? 1979. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (elec-

tronic).  URL http://www.
sciencedirect.com/science/
article/pii/0898122181901413.